# FroM: Frobenius Norm-Based Data-Free Adaptive Model Merging

### v1

@@ Abstract
Zijian Li  Xiaocheng Feng†‡  Huixin Liu Yichong HuangTing LiuBing Qin†‡ Harbin Institute of Technology  Peng Cheng Laboratory {lizijian, xcfeng, hxliu, ychuang, tliu, qinb}@ir.hit.edu.cn
@@ 3.2 Merging of FFT Models
where $n$ denotes the number of fine-tuned models, $\boldsymbol{\theta}_{i}$ represents the task vector of the $i$-th model, $\boldsymbol{\theta}$ is the task vector after merging, and $k$ is a hyperparameter. The comparison between the RegMean and FroM Methods is illustrated in Figure 2.
@@ 3.3 Merging of LoRA
Since Equation 6 is a non-convex function, it may have multiple local minima. Therefore, an alternating optimization algorithm is required to solve it. Specifically, by initializing $\mathbf{A}$ from a normal distribution with mean 00 and variance $\sigma^{2}$, and treating either $\mathbf{A}$ or $\mathbf{B}$ as an invariant, we iteratively update $\mathbf{B}$ and $\mathbf{A}$ using Equation 7 and Equation 8, respectively. In these equations, $\dagger$ denotes the Moore-Penrose inverse (see Appendix A for the proof). This procedure for approximating a local minimum of the objective function is shown in Algorithm 1. Due to the potential computational errors in the calculation of the MP inverse, an early stopping mechanism is implemented, which halts the iteration when the loss begins to increase.
@@ 5.1 Merging of FFT Models
Models Merging Methods
@@
LM & Math Task Arithmetic 50.55 41.62 6.56 / / 32.91
@@
LM & Code Task Arithmetic 53.23 / / 32.93 30.60 38.92
@@
Math & Code Task Arithmetic / 58.45 12.32 5.49 7.00 20.82
@@
LM & Math & Code Task Arithmetic 52.84 42.00 6.20 15.85 18.20 27.02
@@
Models Merging Methods
@@
LM & Math Task Arithmetic 2.25 18.95 6.04 / / 9.08
@@
Math & Code Task Arithmetic / 0.38 0.00 0.00 0.00 0.09
@@
LM & Math & Code Task Arithmetic 0.00 0.45 0.00 0.00 0.00 0.09
@@
For the base model Qwen2.5-7B, we present the results of merging the models Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct, and Qwen2.5-Coder-7B-Instruct, as shown in Table 2. Unexpectedly, the performance of the merged models deteriorates significantly, except when merging Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct using the Task Arithmetic or FroM methods, which yield comparatively better results. When the other two models are integrated with Qwen2.5-Math-7B-Instruct, a substantial decline in performance is observed. Based on the specific test outputs, we observe that the majority of the model’s responses are either incoherent or irrelevant to the given task. This suggests a significant degree of task interference between Qwen2.5-Math-7B-Instruct and the other two models. After extensive fine-tuning on different datasets, the linear mode connectivity Ainsworth et al. (2022); Zhou et al. (2023) among these models tends to be disrupted, leading to the failure of linear weighting methods. This finding indicates that not all fine-tuned models derived from the same base model are equally suitable for model merging. Furthermore, the merging results using the TIES-Merging and DARE methods are close to zero (with outputs consisting entirely of garbled text), indicating that these methods aggravate task interference. In contrast, our FroM method demonstrates greater robustness when merging both two and three models.
@@ 5.2 Merging of LoRA
For the LoRA merging experiments, we train a shared classification head on the MNLI dataset. The performance of this classification head on other datasets is presented in the first row of Table 3. The accuracy across different tasks exceeds 50%, demonstrating a reasonable degree of generalization, considering that for binary classification tasks, the probability of randomly selecting the correct answer is 50%. We further fine-tune on the QNLI, RTE, and WNLI datasets, selecting the best-performing checkpoints for subsequent experiments.
The testing results for different LoRA merging methods are summarized in Table 3. It is evident that the application of the TIES-Merging and DARE methods results in a decline in model accuracy, indicating that these approaches may not be universally effective across all downstream tasks and may perform particularly poorly on simpler ones. On the other hand, our FroM method continues to demonstrate effectiveness, with its average performance surpassing other baselines. Notably, for the QNLI task, our method even exceeds the baseline accuracy, highlighting that our approach enhances the generalization capability of the merged model resulting from multi-task learning.
@@ B.1 Fine-tuning Settings
For the integration of LoRA adapters, the hyperparameters are set as follows: the ranks of matrices A and B are fixed at 16, the batch size is set to 32, and the learning rate is set to $51e-$ 5. We first train a classification head using the MNLI dataset. Since MNLI is a three-class task, we modify it to a binary classification output by combining the neutral and contradiction categories into a single not_entailment class. For each task within the QNLI, RTE, and WNLI datasets, we train the models for 10 epochs and select the checkpoint with the highest accuracy for subsequent LoRA adapter integration.

### v2

@@ Abstract
Zijian Li  Xiaocheng Feng†‡thanks: Corresponding authors.  Huixin Liu Yichong HuangTing LiuBing Qin†‡11footnotemark: 1 Harbin Institute of Technology  Peng Cheng Laboratory {lizijian, xcfeng, hxliu, ychuang, tliu, qinb}@ir.hit.edu.cn
@@ 3.2 Merging of FFT Models
where $n$ denotes the number of fine-tuned models, $\boldsymbol{\theta}_{i}$ represents the task vector of the $i$-th model, $\boldsymbol{\theta}$ is the task vector after merging, and $k$ is a hyperparameter. When $k=0$, the objective function of FroM serves as an upper bound for that of RegMean: $\sum_{i=1}^{n}\|\boldsymbol{W}\boldsymbol{X}-\boldsymbol{W}_{i}\boldsymbol{X}\|_{F}^{2}\leq\sum_{i=1}^{n}\|\boldsymbol{W}-\boldsymbol{W}_{i}\|_{F}^{2}\cdot\|\boldsymbol{X}\|_{F}^{2}=\sum_{i=1}^{n}\|\boldsymbol{\theta}-\boldsymbol{\theta}_{i}\|_{F}^{2}\cdot\|\boldsymbol{X}\|_{F}^{2}$, where $\|\boldsymbol{X}\|_{F}^{2}$ acts as a scaling factor. Moreover, by setting $k>0$, FroM enables more flexible control in determining an appropriate trade-off point. The comparison between the RegMean and FroM Methods is illustrated in Figure 2.
@@ 3.3 Merging of LoRA
Since Equation 6 is a non-convex function, it may have multiple local minima. Therefore, inspired by the Alternating Direction Method of Multipliers in convex optimization theory and analogous to LoRM Salami et al. (2024), we propose an alternating optimization algorithm. Specifically, by initializing $\mathbf{A}$ from a normal distribution with mean 0 and variance $\sigma^{2}$, and treating either $\mathbf{A}$ or $\mathbf{B}$ as an invariant, we iteratively update $\mathbf{B}$ and $\mathbf{A}$ using Equation 7 and Equation 8, respectively. In these equations, $\dagger$ denotes the Moore-Penrose inverse (see Appendix A for the proof). This procedure for approximating a local minimum of the objective function is shown in Algorithm 1. Due to the potential computational errors in the calculation of the MP inverse, an early stopping mechanism is implemented, which halts the iteration when the loss begins to increase.
@@ 5.1 Merging of FFT Models
Models
Merging
Methods
@@
LM
& Math
Task Arithmetic 50.55 41.62 6.56 / / 32.91
@@
LM
& Code
Task Arithmetic 53.23 / / 32.93 30.60 38.92
@@
Math
& Code
Task Arithmetic / 58.45 12.32 5.49 7.00 20.82
@@
LM
& Math
& Code
Task Arithmetic 52.84 42.00 6.20 15.85 18.20 27.02
@@
Models
Merging
Methods
@@
LM
& Math
Task Arithmetic 2.25 18.95 6.04 / / 9.08
@@
Math
& Code
Task Arithmetic / 0.38 0.00 0.00 0.00 0.09
@@
LM
& Math
& Code
Task Arithmetic 0.00 0.45 0.00 0.00 0.00 0.09
@@
For the base model Qwen2.5-7B, we present the results of merging the models Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct, and Qwen2.5-Coder-7B-Instruct, as shown in Table 2. Unexpectedly, the performance of the merged models deteriorates significantly, except when merging Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct using the Task Arithmetic or FroM methods, which yield comparatively better results. When the other two models are integrated with Qwen2.5-Math-7B-Instruct, a substantial decline in performance is observed. Based on the specific test outputs, we observe that the majority of the model’s responses are either incoherent or irrelevant to the given task. This suggests a significant degree of task interference between Qwen2.5-Math-7B-Instruct and the other two models. After extensive fine-tuning on different datasets, the linear mode connectivity Ainsworth et al. (2022); Zhou et al. (2023b) among these models tends to be disrupted, leading to the failure of linear weighting methods. This finding indicates that not all fine-tuned models derived from the same base model are equally suitable for model merging. Furthermore, the merging results using the TIES-Merging and DARE methods are close to zero (with outputs consisting entirely of garbled text), indicating that these methods aggravate task interference. In contrast, our FroM method demonstrates greater robustness when merging both two and three models.
Since the training data of these FFT models are not fully released, it is impossible to directly compare the FroM and RegMean methods in Table 1 and Table 2. That said, we fully agree that including RegMean as a baseline is crucial for a more comprehensive comparison. To this end, we conduct additional experiments with RegMean, with detailed results presented in Appendix C.
@@ 5.2 Merging of LoRA
For the LoRA merging experiments, we train a shared classification head on the MNLI dataset. The performance of this classification head on other datasets is presented in the first row of Table 5.2. The accuracy across different tasks exceeds 50%, demonstrating a reasonable degree of generalization, considering that for binary classification tasks, the probability of randomly selecting the correct answer is 50%. We further fine-tune on the QNLI, RTE, and WNLI datasets, selecting the best-performing checkpoints for subsequent experiments.
The testing results for different LoRA merging methods are summarized in Table 5.2. It is evident that the application of the TIES-Merging and DARE methods results in a decline in model accuracy, indicating that these approaches may not be universally effective across all downstream tasks and may perform particularly poorly on simpler ones. On the other hand, our FroM method continues to demonstrate effectiveness, with its average performance surpassing other baselines. Notably, for the QNLI task, our method even exceeds the baseline accuracy, highlighting that our approach enhances the generalization capability of the merged model resulting from multi-task learning.
@@ 6 Conclusion
Acknowledgments
Xiaocheng Feng and Bing Qin are the co-corresponding authors of this work. We thank the anonymous reviewers for their insightful comments. This work was supported by the National Natural Science Foundation of China (NSFC) (grant 62276078, U22B2059), the Key R&D Program of Heilongjiang via grant 2022ZX01A32, and the Fundamental Research Funds for the Central Universities (XNJKKGYDJ2024013).
@@ B.1 Fine-tuning Settings
For the integration of LoRA adapters, the hyperparameters are set as follows: the ranks of matrices A and B are fixed at 16, the batch size is set to 32, and the learning rate is set to 1e51e-5. We first train a classification head using the MNLI dataset. Since MNLI is a three-class task, we modify it to a binary classification output by combining the neutral and contradiction categories into a single not_entailment class. For each task within the QNLI, RTE, and WNLI datasets, we train the models for 10 epochs and select the checkpoint with the highest accuracy for subsequent LoRA adapter integration.
@@ B.2 Test Details
Appendix C Additional Comparison with RegMean
We evaluate scenarios involving the merging of three and four models, in order to better demonstrate the generalization capability of our FroM method. Specifically, we fine-tune the Meta-Llama-3-8B base model on four tasks using supervised fine-tuning: Instruction Following, Mathematical Reasoning, Code Generation, and Safety. The corresponding datasets are Alpaca-cleaned Taori et al. (2023), NuminaMath-CoT LI et al. (2024), Code Alpaca Chaudhary (2023), and Saferpaca Bianchi et al. (2023). For evaluation across these four dimensions, we use IFEval Zhou et al. (2023a), GSM8K, HumanEval, and DirectHarm4 Lyu et al. (2024), where the metric is the proportion of safe responses. We merge three and four models to better evaluate the generalization ability of these methods. The results are shown in Table C. Since RegMean leverages a large amount of training data, FroM does not outperform RegMean but still achieves better results than the other methods.
The results of the linear search for the linear weighting coefficient and the hyperparameter used in FroM within the range $[0.5,1.0]$ are presented in Table 8 and Table 9. The final choices are $\alpha=0.6$ and $k=0.5$.
Merging Methods IFEval GSM8K HumanEval DirectHarm4 Avg.
Baseline
/ 21.34 38.36 35.37 57.75 38.21
Fine-tuning
/ 41.13 42.76 45.73 93.00 55.66
3-Model Merging
Task Arithmetic 44.48 48.29 44.51 / 45.76
TIES-Merging 44.48 39.42 43.90 / 42.60
DARE 26.38 0.00 0.00 / 8.79
DARE+TIES-Merging 26.38 0.00 0.00 / 8.79
RegMean 43.65 52.08 45.12 / 46.95
[4pt/5pt] FroM 44.48 49.89 46.34 / 46.90
DARE+FroM 37.05 33.74 2.44 / 24.41
4-Model Merging
Task Arithmetic 43.17 51.55 36.59 71.25 50.64
TIES-Merging 43.65 44.73 45.73 64.50 49.65
DARE 26.38 0.00 0.00 70.50 24.22
DARE+TIES-Merging 26.38 0.00 0.00 67.25 23.41
RegMean 43.17 54.97 42.07 97.75 59.49
[4pt/5pt] FroM 45.20 52.01 45.12 65.25 51.90
DARE+FroM 21.94 32.68 0.61 72.50 31.93
Table 7: Merging results of three and four models based on Meta-LLaMA-3-8B, in comparison with RegMean.
$\alpha$ IFEval GSM8K HumanEval DirectHarm4 Avg.
3-Model Merging
$0.5$ 43.88 49.96 43.90 / 45.91
$0.6$ 44.48 48.29 44.51 / 45.76
$0.7$ 41.25 47.38 32.93 / 40.52
$0.8$ 41.61 41.77 20.73 / 34.70
$0.9$ 40.41 37.83 9.15 / 29.13
$1.0$ 37.17 33.36 2.44 / 24.32
4-Model Merging
$0.5$ 43.65 52.08 42.07 64.25 50.51
$0.6$ 43.17 51.55 36.59 71.25 50.64
$0.7$ 39.21 49.51 23.17 74.50 46.60
$0.8$ 39.57 46.47 7.93 73.50 41.87
$0.9$ 29.14 43.06 0.61 64.75 34.39
$1.0$ 21.94 32.68 0.61 72.75 32.00
Table 8: Results of the linear search for the weighting coefficient used for additional comparison with RegMean.
$k$ IFEval GSM8K HumanEval DirectHarm4 Avg.
3-Model Merging
$0.5$ 44.48 49.89 46.34 / 46.90
$0.6$ 44.48 48.07 45.73 / 46.09
$0.7$ 43.41 47.01 47.56 / 45.99
$0.8$ 43.88 48.67 45.73 / 46.09
$0.9$ 43.17 47.76 45.73 / 45.55
$1.0$ 43.76 46.25 43.90 / 44.64
4-Model Merging
$0.5$ 45.20 52.01 45.12 65.25 51.90
$0.6$ 45.44 49.66 45.73 64.25 51.27
$0.7$ 46.04 48.52 46.95 64.00 51.38
$0.8$ 44.00 46.40 46.34 69.50 51.56
$0.9$ 43.41 47.99 48.17 66.75 51.58
$1.0$ 42.45 46.40 44.51 65.00 49.59
Table 9: Results of the linear search for $k$ used for additional comparison with RegMean.